# R+X: Retrieval and Execution from Everyday Human Videos

### v1

@@ 3.2 Execution: Few-Shot, In-Context Imitation from Video Examples
Given the output of the retrieval phase $[\Lambda,\mathcal{M}]$ and a new, live RGB-D observation collected by the robot $\mathcal{O}_{live}$, we extract its visual 3D keypoints representation $\mathcal{K}_{live}$ by first extracting its $N\times 384$ DINO descriptors, and then finding the $K$ nearest neighbours to each of the $K$ descriptors in $\mathcal{D}$, that we obtained in the retrieval phase. This results in $K$ 2D coordinates, that we project in 3D. We then input to KAT as context $[\Lambda,\mathcal{M}]$, as $Z$ examples of the desired input-output mapping, and the new visual 3D keypoints, and generate a new trajectory of desired hand joint actions, $\text{KAT}([\Lambda,\mathcal{M}],\mathcal{K}_{live})\rightarrow\mathcal{J}_{live}$ with $\mathcal{J}_{live}=[j_{1},\dots,j_{T}]$, as shown in Figure 4. We then map the predicted trajectories of hand joints actions into gripper poses as described in our Supplementary Material, and execute them.
To summarize, given a language command $\mathcal{L}$ and a human video $\mathcal{H}$, R+X first retrieves $Z$ videos depicting the described behaviour using a VLM. From the retrieved videos, a list of $[\Lambda,\mathcal{M}]$ visual 3D points and hand joint actions are extracted. Then, to execute the desired behaviour described in the prompt $\mathcal{L}$, $[\Lambda,\mathcal{M}]$ along with the visual 3D keypoints of the live observation $\mathcal{K}_{live}$ are used as context for KAT, which performs few-shot in-context imitation learning to generate a trajectory of hand joints actions, which is then mapped to gripper poses and executed on the robot.
@@
4 Experiments
Human Video. We collect the human video $\mathcal{H}$ using an Intel RealSense 455, worn by the human on their chest as shown in Figure 1. To reduce downstream computational time, we filter out each frame in which human hands are not visible right after recording. As our robot is single-armed, we limit ourselves to single-hand tasks. However, our method could be applied to bimanual settings and dexterous manipulators.
Robot Setup. At execution, we use a Sawyer robot equipped with a RealSense 415 head-camera and a two-fingered parallel gripper, the Robotiq 2F-85. As the robot is not mobile, we set up different scenes in front of it with variations of the tasks recorded by the human, placing several different distractors for each task, while the human video was recorded in many different rooms. Although we have a wrist-camera mounted to the robot as shown in some of our figures, we do not use it in our work.
Tasks. To evaluate our proposed framework, we use a set of 12 everyday tasks, where the human interacts with a series of common household objects, listed in Fig. 6. We include movements like grasping, opening, inserting, pushing, pressing, and wiping.
@@ 4 Experiments
More details on all aspects are provided in the Supplementary Material.
@@
Table 1: Success rates of the various methods on the 12 proposed tasks.
4.1 Results
@@
Can R+X learn robot skills from long, unlabelled videos? How does it perform with respect to a monolithic language-conditioned policy? In these experiments, we evaluate the performance of R+X in learning a set of everyday skills. At deployment, we place the object to be interacted with in front of the robot, and issue a language command. We then run the retrieval and execution pipeline using the human video, the current observation and the language command. We run 10 episodes per task, for each episode randomising 1) the object pose, 2) the type and number of distractors, 3) for tasks where it is possible, we swap the object to interact with, with another one from the same class to test for generalisation (e.g. a different can, a different piece of clothing, a different telephone). More details are provided in the Supplementary Material.
Figure 7: Examples of hard spatial and language generalisation. Gripper moves from red to blue.
@@
In Table 1, we report the performance of R+X on these tasks, together with the baselines. As we demonstrate, the framework is able to tackle a wide range of everyday tasks, surpassing the monolithic policy baselines. These results prove the benefit of modeling language-conditioned learning from observation as distinct retrieval and execution phases, to fully leverage the abilities of recent Foundation Models.
@@
What are the main sources of difference in performance between R+X and a monolithic policy? In these experiments, we investigate more in detail what changes in the inputs lead to the most noticeable difference in performance between R+X and the baselines, R3M-DiffLang and Octo. We explored two aspects, related to two properties of R+X:
@@
Hard Spatial Generalisation: R+X, by retrieving videos of the desired task, can also extract a series of relevant keypoints $\mathcal{K}_{live}$ from the current observation $\mathcal{O}_{live}$, something not possible when using a single policy network. Going from RGB-D to a list of 3D points for inputs and outputs allows us to apply simple geometric data augmentation techniques for KAT, such as normalisation or random translations and rotations. This leads to a stronger spatial generalisation: in the "grasp a can" and "grasp a beer" tasks, we test performance of each method when the objects are on the table, or when they are positioned on top of other objects (a box, a microwave), a situation never encountered in the training videos. By running 5 test episodes for each case, we demonstrate how R+X retains strong performance, while the performance of the baselines drop. Results and an example of the predicted actions can be seen in Fig. 7.
@@
Hard Language Generalisation: By leveraging the language and video understanding abilities of recent large Vision Language Models, such as Gemini, R+X can interpret and execute nuanced commands. To evaluate this, we set up a scene with many objects, and ask the robot to perform three tasks: "give me something to call my mom", "give me a non-alcoholic drink", and "make the room less dark". We run 5 test episodes for each of these commands, modifying the positions of the objects and the distractors. The language-conditioned policies struggle to interpret these commands, since they are strongly out of distribution. R+X, on the other hand, leverages Gemini’s ability to understand the meaning of these commands, as it is able to retrieve useful video clips from the human video (respectively, picking up the phone, grasping a Fanta can, and turning on the light). Results and an example of output gripper trajectories can be seen in Fig. 7.
@@
Figure 8: Left (barplots 1-3): We compare R+X and the baselines’ ability to learn tasks in succession, and the time needed to learn such new tasks. Rightmost barplot: Gemini’s retrieval performance.
@@
Can R+X learn task sequentially over time? In these experiments, we demonstrate how R+X can learn tasks continually, with no need for any additional training or finetuning, while obtaining strong performance both on the new task and on the old ones, a desirable ability for a robot learning from an ever increasing dataset of human experience. To measure this ability, and highlight the different behaviour with respect to a single language-conditioned policy, we first collect 10 demos for 3 tasks: grasping a can, cloth in basket, and picking up the phone. We train the baselines on these tasks (after extracting captions via Gemini as described for the experiments of Table 1) and evaluate them and R+X. Then, we add 10 demos of 3 new tasks (kettle on stove, open box, plate in rack), finetune the baselines, then measure performance on the 3 new tasks, and on the 3 old tasks, for all methods. In Fig. 8, Left, we see how the performance of R3M-DiffLang and Octo deteriorates on the old tasks, due to the well known effect of catastrophic forgetting [34]. On the other hand, R+X performance does not deteriorate, due to the ability to retrieve from an ever-growing video of tasks and adapting the behaviour model at test time on the retrieved data.
@@
If instead we train the baselines on all the data each time, the performance does not deteriorate: however, this leads to a substantial growth in time needed to train. As the dataset grows, those networks would need an increasing amount of time per each new added task in order to be retrained, while R+X does not need any training or finetuning as further tasks are added to the video.
How does the length of the videos affect the retrieval performance? In this experiment, we first collect a video with 10 demos of 5 tasks (kettle on stove, cloth in basket, pick up phone, cloth in washing machine, close microwave), and measure Gemini’s ability to retrieve clips for each task. We then add 10 demos for 5 additional tasks, and compute again the precision and recall of retrieval of the original 5 tasks, to study how length of video and number of tasks affects it. The original video is close to 4 minutes in length, while the second is close to 10 minutes. In Fig. 8, Right, we show the mean over all tasks of the retrieval’s precision and recall. Results demonstrate how Gemini’s performance remains high while increasing the length of the video, allowing to scale R+X as the human records more and more tasks.
5 Discussion and Limitations
We have presented R+X, a framework to learn robot skills from long, unlabelled videos of humans interacting naturally in everyday environments, and demonstrated the clear benefit of using retrieval over training monolithic language-conditioned policy networks. There are, however, limitations which we now describe, opening up possible avenues for future work. The method is currently bottlenecked by the performance of Vision Language Model; however, such models have seen dramatic improvement in recent times and this will likely continue. While we can learn a large series of everyday tasks, the errors arising from the hand pose prediction currently inhibits us from learning precise tasks. Furthermore, while the keypoints extraction leads to strong spatial generalisation properties, it can become a bottleneck in the presence of many similar objects; methods that find a dynamic number of keypoints might tackle this issue. Finally, due to the use of Foundation Models for retrieval and execution, our method has a few seconds of delay between the issuing of the command and execution; however, this has reduced dramatically in recent months with the release of faster models such as GPT-4o [35] or Gemini Flash [24].

### v2

@@ III-B Execution: Few-Shot, In-Context Imitation from Video Examples
Given the output of the retrieval phase $[\Lambda,\mathcal{M}]$ and a new, live RGB-D observation collected by the robot $\mathcal{O}_{live}$, we extract its visual 3D keypoints representation $\mathcal{K}_{live}$ by first extracting its $N\times 384$ DINO descriptors, and then finding the $K$ nearest neighbours to each of the $K$ descriptors in $\mathcal{D}$, that we obtained in the retrieval phase. This results into $K$ 2D coordinates, that we project in 3D. We then input to KAT as context $[\Lambda,\mathcal{M}]$, as $Z$ examples of the desired input-output mapping, and the new visual 3D keypoints, and generate a new trajectory of desired hand joint actions, $\text{KAT}([\Lambda,\mathcal{M}],\mathcal{K}_{live})\rightarrow\mathcal{J}_{live}$ with $\mathcal{J}_{live}=[j_{1},\dots,j_{T}]$, as shown in Figure 3. We then map the predicted trajectories of hand joints actions into gripper poses as described in our website and appendix, and execute them.
To summarize, given a language command $\mathcal{L}$ and a human video $\mathcal{H}$, R+X first retrieves $Z$ videos depicting the described behaviour using a VLM. From the retrieved videos, a list of $[\Lambda,\mathcal{M}]$ visual 3D points and hand joint actions are extracted. Then, to execute the desired behaviour described in the prompt $\mathcal{L}$, $[\Lambda,\mathcal{M}]$ along with the visual 3D keypoints of the live observation $\mathcal{K}_{live}$ are used as context for KAT that performs few-shot in-context imitation learning to generate a trajectory of hand joints actions, which is mapped to gripper poses and executed on the robot.
IV Experiments
Human Video. We collect the human video $\mathcal{H}$ using an Intel RealSense 455, worn by a human on their chest as shown in Figure R+X: Retrieval and Execution from Everyday Human Videos. To reduce downstream computational time, we filter out each frame in which human hands are not visible right after recording. As our robot is single-armed, we limit ourselves to single hand tasks. However, our method could identically be applied to bimanual settings and dexterous manipulators. The video is collected in many different rooms and buildings.
Robot Setup. At execution, we use a Sawyer robot equipped with a RealSense 415 head-camera. The robot is equipped with a two-fingered parallel gripper, the Robotiq 2F-85. As the robot is not mobile, we setup different scenes in front of it with variations of the tasks recorded by the human, placing several different distractors for each task, while the human video was recorded in many different rooms. Although we have a wrist-camera mounted, we do not use that in our work.
Tasks. To evaluate our proposed framework, we use a set of 12 everyday tasks, where a human interacts with a series of common household objects, listed in Fig. 4. We include movements like grasping, opening, inserting, pushing, pressing, and wiping.
@@ IV Experiments
Videos, code, and additional implementation details can be found on our website: https://www.robot-learning.uk/r-plus-x
@@
TABLE I: Result of the various methods on the 12 proposed tasks.
IV-A Results
@@
Can R+X learn robot skills from long, unlabelled videos? How does it perform with respect to a monolithic language-conditioned policy? In these experiments, we evaluate the performance of R+X in learning a set of everyday skills. At deployment, we place the object to be interacted with in front of the robot, and issue a language command. We then run the retrieval and execution pipeline using the human video, the current observation and the language command. We run 10 evaluations per task, randomising 1) the object pose 2) the type and number of distractors 3) for tasks where it is possible, we swap the object to interact with, with another one from the same class to test for generalisation (e.g. a different can, a different piece of clothing, a different telephone). We provide more details in the appendix.
@@
In Table I, we report the performance of R+X on these tasks, together with the baselines. As we demonstrate, the framework is able to tackle a wide range of everyday tasks, surpassing the monolithic policy baselines. These results prove the benefit of modeling language-conditioned learning from observation as distinct retrieval and execution phases, to fully leverage the abilities of recent Foundation Models. In the appendix, we investigate more in depth the performance of R+X on unseen objects and in the presence of distractors.
Figure 5: Examples of spatial, language and distractors generalisation. Gripper trajectories move from red to blue.
@@
What are the main sources of difference in performance between R+X and a monolithic policy? In these experiments, we investigate what changes in the inputs lead to the most noticeable difference in performance between R+X and the baselines, R3M-DiffLang and Octo. We explored two aspects, related to two properties of R+X:
@@
Hard Spatial Generalisation: R+X, by retrieving videos of the desired task, can also extract a series of relevant keypoints $\mathcal{K}_{live}$ from the current observation $\mathcal{O}_{live}$, something not possible when using a single policy network. Going from RGB-D to a list of 3D points for inputs and outputs allows us to apply simple geometric data augmentation techniques for KAT, such as normalisation or random translations and rotations. This leads to stronger spatial generalisation: in the ”grasp a can” and ”grasp a beer” tasks we test each method’s performance when the objects are on the table, or when they are positioned on top of other objects (a box, a microwave). By running 5 test rollouts for each case, we demonstrate how R+X retains strong performance, while the baselines’ performance drops. Results and an example of the predicted gripper trajectories can be seen in Fig. 5, middle.
@@
Hard Language Generalisation: By leveraging the language and video understanding abilities of recent large Vision Language Models, such as Gemini, R+X can interpret and execute nuanced commands. To evaluate this, we setup a scene with many objects, and ask the robot to perform three tasks: ”give me something to call my mom”, ”give me a non-alcoholic drink” and ”make the room less dark”. We run 5 test rollouts for each of these commands, modifying the position of the objects and the distractors. The language-conditioned policies struggle to interpret these commands, that are strongly out of distribution. R+X, on the other hand, leverages Gemini’s ability to understand the meaning of these commands, as it is able to retrieve useful video clips from the human video (respectively, picking up the phone, grasping a Fanta can, and turning on the light). Results and an example of output gripper trajectories can be seen in Fig. 5, left.
@@
Figure 6: We compare R+X and the baselines’ ability to learn tasks in succession, and the time needed to learn such new tasks.
Distractors Generalisation: One of the main differences between R+X and the baselines is that the former, after retrieving a set of video clips from $\mathcal{H}$ depicting the requested task, can extract a set of keypoints that are semantically and geometrically meaningful for the objects to interact with, as described in the main paper. These keypoints are generally robust to distractors, as they focus only on the DINO features that are common in all images, whereas distractors typically vary throughout the human video $\mathcal{H}$ and the execution. The baselines, being monolithic policies, receive as input the observation as it is, and must learn to generalise to the presence of distractors during the long training phase.
To evaluate the effect of distractors on the scene, we test R+X and the baselines on two tasks, ”grasp a can” and ”put the cloth in the basket”, emulating the experimental scenario of the ”Hard Generalisations” sections. We run 10 runs without distractors, and 10 with, explicitly measuring the performance in the two cases. We can see in the results of Fig. 5 right, how R+X is more robust to the presence of random distractor objects, highlighting the advantage of extracting semantically and geometrically meaningful keypoints after having received a language command and having retrieved the corresponding video clips.
@@
Can R+X learn task sequentially over time? In these experiments, we demonstrate how R+X can learn tasks continually, with no need for any additional training or finetuning, while obtaining strong performance both on the new task and on the old ones, a desirable ability for a robot learning from an ever increasing dataset of human experience. To measure this ability, and highlight the difference behaviour with respect to a single language-conditioned policy, we first collect 10 demos for 3 tasks. We train the baselines on these tasks (after extracting captions via Gemini as described for the experiments of Table I) and evaluate them and R+X. Then, we add 10 demos of 3 new tasks, finetune the baselines, then measure performance on the 3 new tasks, and on the 3 old tasks for all methods. In Fig. 6, we see how the performance of R3M-DiffLang and Octo deteriorates on the old task, due to the well known effect of catastrophic forgetting [32]. On the other hand, R+X performance does not deteriorate, due to the ability to retrieve from an ever-growing video of tasks and adapting the behaviour model at test time on the retrieved data.
@@
If instead we train the baselines on all the data each time, the performance does not deteriorate: however, this leads to a substantial growth in time needed to train. As the dataset grows, those networks would need an increasing amount of time per each new added task in order to be retrained, while R+X does not need any training or finetuning.
V Conclusion
We presented R+X, a method to learn robot skills from long, unlabelled videos of humans interacting with their environments. Given a language command and a live observation of the scene, R+X retrieves a sequence of relevant video clips from a long unlabelled video which it uses as context for an in-context imitation learning method that predicts what actions to execute. Through several real-world experiments on 12 everyday tasks, we demonstrate that R+X can successfully generalise to previously unseen objects, scenes, and object placements, is robust to distractor objects, and its use of VLMs for video retrieval enables R+X to generalize to complex language commands demonstrating its clear benefits over training monolithic language-conditioned policy networks. For a discussion on limitations and future work please see the appendix.